OpenStudy (anonymous):

Find the volume of this composite solid. A. 1099.01 m3 B. 785 m3 C. 104.67 m3 D. 889.67 m3

OpenStudy (anonymous):



OpenStudy (anonymous):

*10m PLEASE HELP i need answer ASAP

OpenStudy (mathmate):

Hints: \( Volume\ of\ cylinder = \pi r^2 h \\ Volume\ of\ cone = \pi r^2h/3 \) h refers to the corresponding height (of cylinder or cone).

OpenStudy (anonymous):

soo cylinder = 314? && the cone = 261.66~ ???

OpenStudy (mathmate):

Not quite, can you please show your work?

OpenStudy (mathmate):

Watch out that \( r^2 \) means r*r .

OpenStudy (anonymous):

cylinder: 3.14(25)4 cone: 3.14(25)(10)/3 cylinder=314 cone=261.66~ thats what i got ?

OpenStudy (mathmate):

The heights of the cylinder and cone have been interchanged. I don't know if that's a mistake or a wrong diagram. Also, the formulae call for \( r^2 \) which means r*r.

OpenStudy (anonymous):

if i put the link of the picture would you be able to see it?

OpenStudy (mathmate):

I would, but not necessary as long as you have checked that the numbers are correct. Your picture shows a cylinder of 10 high and a cone 4 high, but you have interchanged them in your calculations, hence my doubt.

OpenStudy (anonymous):

ive checked and they are correct

OpenStudy (anonymous):

@mathstudent55 can you please help me?

OpenStudy (mathmate):

Did you know that the heights you used in your calculations do not correspond to those given in the diagram, which you said are correct?